Circular No. 8036

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

EDITORIAL NOTICE All contributors of information for IAUCs should note that any such

communication must be sent to CBAT@CFA.HARVARD.EDU, as the Director’s personal e-mail may not always be monitored in his absence.

The Central Bureau has established a new communication outlet, the ‘Electronic Telegrams’ (or CBETs), as a temporary means for getting urgent items to the astronomical community prior to the formal issuance of those items on these Circulars (which are editorially more complex and time-intensive to produce). The intention is for the CBETs to be patterned very loosely after the Telegrams that were issued by the Central Bureau from the 1920s until the 1990s (except that the CBETs will not be in code!). CBETs will initially appear only as e-mail communications to Computer-Service subscribers; it is anticipated that these CBETs will be issued less frequently than IAUCs, and probably chiefly at times when the Director is unavailable. Line-charge items contributed for these IAUCs may occasionally appear first on the CBETs (for urgent items), but only one charge will be assessed.

V838 MONOCEROTIS D. P. K. Banerjee and N. M. Ashok, Physical Research Laboratory

(PRL), Ahmedabad, report J-band imaging and near-infrared spectroscopy at 1.08–2.35 μm, with the Mt. Abu 1.2-m telescope (+ PRL Near Infrared Imaging Spectrometer + HgCdTe NICMOS3 detector array) of V838 Mon. J-band images taken on Dec. 14 clearly detect a faint shell around V838 Mon (seen earlier only in the optical). The shell (diameter ≈ 60′′) is fairly spherical but more extended in the east-west direction. The brightest Jband feature is an enhanced bar structure seen in the north-west direction. JHK spectroscopy (resolution ∼ 1000) confirms earlier reports (IAUC 8016) that the near-infrared spectrum of V838 Mon has evolved significantly. The previously reported absorption bands at 1.23, 1.25, 1.65, and 1.69 μm and the first-overtone CO bands have persisted and are prominently seen during three recent observations (Nov. 20, Dec. 7, 13). The H-band absorption features are especially strong, while the CO band structure is rather similar to its appearance during the second outburst of V838 Mon on 2002 Feb. 2 (Banerjee and Ashok 2002, A.Ap. 395, 161).

2002 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green